How many joules of heat are absorbed when 70.0 grams of water is completely vaporized at its boiling point?


1
Expert's answer
2021-01-18T01:54:57-0500

water enthalpy of vaporization is 40.65 kJ/mol

n(H2O) = (70 g)/(18 g/mol) = 3.8889 mol

"\Delta"H = 3.8889*40.65 = 158.0833 kJ = 158083.3 J